Two blocks are fastened to the ceiling of an elevator. The elevator accelerates upward at 1.50 m/s2. The blocks both have a mass of 12.5 kg. Find the tension in each rope.

top rope N

bottom rope N
Answer #1

uploaded image

Tension in the lower one = m(g+a) = 12.5*(9.8+1.50) = 141.25 N


Tension in the upper one = 2m(g+a) = 2*12.5*(9.8+1.50) = 282.5 N